# The causal structure of galactic astrophysics

### v2

@@ 3.3 Mock data generation
This is achieved with the Causal Perceptron Network (CPN; Ramsey et al. 2025b), a simulation framework for generating synthetic datasets from arbitrary nonlinear models. The user specifies a DAG that encodes the desired causal structure, along with a noise distribution. Each variable is then expressed as a nonlinear function of its causes plus an independent noise term. Rather than choosing simple algebraic forms for these functions CPN uses randomly configured multilayer perceptrons, ensuring that the resulting data exhibits realistic nonlinear dependencies while remaining stable even with a large number of variables. Each dataset is made by recursively sampling noise and propagating values forward through the causal graph, producing independent and identically distributed samples. This lets CPN produce highly flexible data approximating a broad class of nonlinear functions.
We generate datasets with the same size as the NSA subset (587,338 points) with 7 nodes and a random number of edges between 12 and 16, roughly matching what will be measured in the real data. This takes around a minute per dataset. We use four hidden layers with 50 neurons each, a ReLU activation function and the default noise distribution $B(2,5)$. This produces mock datasets with correlations visually similar to Fig. 1, but we also check that the results are not sensitive to reasonable variations in them. We then refit each of these datasets with the FCIT algorithm for a range of truncation_limit and penalty_discount values. For each one we compute the precision (fraction of predicted edges that are correct), recall (fraction of true edges that were successfully recovered) and F1 score (harmonic mean of the precision and recall) of the PAG produced.
@@ 4.1 Mock data tests
The result is shown in Fig. 2. As penalty_discount increases the score penalises model complexity more strongly, leading to sparser graphs. This reduces false positives and thus tends to increase precision, but also causes some true edges to be missed, lowering recall. The F1 score peaks at penalty_discount $\sim 40-50$ at a value $\sim 0.9$, roughly indicating a 90 per cent success rate on each dataset. We adopt a value of 50 for the real data, but again explore reasonable variations without finding important differences. We also confirm that the other hyperparameters in the FCIT algorithm and its testing and scoring methods impact the results little.
@@ 4.2 Real data
The result describes a combination of physical effects carrying information about galaxy evolution and observational and selection effects describing the way in which the data was obtained. As expected, redshift influences the apparent size, which scales inversely with angular diameter distance. It also influences mass and absolute magnitude through Malmquist bias, the preferential detection of intrinsically brighter objects at higher distance. Mass is seen to causally determine size (rather than vice versa, as would be possible given simply the mass–size relation), suggesting inside-out growth of discs and size expansion via mergers. It also determines Sérsic index, agreeing with the idea that bulge growth and morphological transformation are primarily consequences of hierarchical mass assembly. The absence of a link from morphology to mass disfavours simplistic models where concentration alone sets stellar mass. The link from SFR to absolute magnitude reflects the brightening of galaxies in optical bands due to recent star formation.
The uncertain edges between star formation, stellar mass and morphology highlight the complexity of baryonic processes. Several appear with circle endpoints, indicating algorithmic uncertainty about direction or the influence of confounding variables. The ambiguous link between stellar mass and luminosity is unsurprising, since mass estimates are derived from photometry and strongly depend on mass-to-light ratios. The circle endpoints highlight the possible role of latent factors not included in the analysis—such as dust attenuation, gas content, and halo environment—which can drive correlations and obscure true directions. The graph does not unambiguously support a picture in which star formation determines morphology on short timescales, or that mass quenching is the sole pathway. Instead, it suggests an intertwined system in which stellar mass remains the fundamental regulator, while star formation and morphology are mutually influenced by common drivers such as gas inflow and environment. Disentangling these physical drivers from observational and selection-induced structure (and external latent variables) requires further work, as indicated directly by the circle edges. The result does however imply that the backbone of galaxy evolution—mass driving size and morphology, and star formation driving luminosity—is recoverable directly from survey data. This is highly promising for future, more sophisticated applications of the methodology.
@@ 5 Discussion and Conclusion
To illustrate the approach we have applied causal discovery to low-redshift galaxy data from the NSA, adapting a hybrid constraint-and-score algorithm—FCIT—to meet the demands of astrophysical data (large datasets, highly nonlinear correlations and presence of confounders). After testing and calibrating the method on NSA-like mock data (establishing $\sim$90 per cent accuracy) we applied it to the real data to find the PAG in Fig. 3. This supports a hierarchical and mass-driven framework of galaxy evolution while indicating the complexities involved in the physical mechanisms at play. It also highlights the vital importance of observational causal discovery methods, since intervention is impossible in astrophysics. Such methods are sometimes undervalued in favour of intervention-based methods involving randomised control trials.
@@
General and efficient causal discovery methods remain under active development. Future algorithms might be able to resolve circle endpoints to distinguish models within a PAG equivalence class by utilising other types of information (e.g. nonlinearity, non-Gaussianity or heteroskedasticity), as has already been done when assuming causal sufficiency in the linear, non-Gaussian regime (e.g. Shimizu et al. 2011). One could even do Bayesian model comparison between competing simulations or theories based on their causal structures they predict (Dhir et al., 2023; Jin et al., 2025b), directly demonstrating the gain in constraining power afforded by causal discovery. Taken together, these steps would provide a more complete and reliable causal map, ultimately allowing causal discovery to move beyond recovering broad backbones to testing detailed models of how galaxies grow and evolve.
@@
The Nasa Sloan Atlas is publicly available at https://www.sdss4.org/dr17/manga/manga-target-selection/nsa/. The code used in this paper will be made publicly available on GitHub upon publication.
@@ Acknowledgements
We thank David Bacon, Deaglan Bartlett, Robin Evans, Pedro Ferreira, Sebastian von Hausegger, Matt Jarvis, Richard Stiskalek and Tariq Yasin for useful discussions. HD is supported by a Royal Society University Research Fellowship (grant no. 211046).

### v3

@@ 3.3 Mock data generation
This is achieved with the Causal Perceptron Network (CPN; Ramsey et al. 2025b), a simulation framework for generating synthetic datasets from arbitrary nonlinear models. The user specifies a DAG that encodes the desired causal structure, along with a noise distribution. Each variable is then expressed as a nonlinear function of its causes plus an independent noise term. Rather than choosing simple algebraic forms for these functions CPN uses randomly configured multilayer perceptrons, ensuring that the resulting data exhibits realistic nonlinear dependencies while remaining stable even with a large number of variables.
Each dataset is made by recursively sampling noise and propagating values forward through the causal graph, producing independent and identically distributed samples. In detail, this means that for each variable $X_{i}$, a structural equation of the form
$X_{i}=f_{i}\!\left(\mathrm{Pa}(X_{i}),\,E_{i}\right)$ (3)
is defined, where $\mathrm{Pa}(X_{i})$ denotes the set of parent variables of $X_{i}$ in the DAG, $f_{i}$ is a feed-forward neural network, and $E_{i}$ is an independent exogenous noise term. The noise variables $E_{i}$ are sampled independently and then the variables computed in the causal order of the DAG. This forward propagation ensures that each variable depends only on its parents and its own noise term. Because the noise enters as an explicit input to the neural network rather than being added post hoc, the resulting conditional distributions $p(X_{i}\mid\mathrm{Pa}(X_{i}))$ can be highly nonlinear, non-Gaussian, and heteroskedastic. This lets CPN produce highly flexible data suitable for testing causal discovery methods under realistic nonlinear data-generating processes. Note that the measured values from the real dataset are not used in this procedure.
We generate datasets with the same size as the NSA subset (445,763 points) with 7 nodes and a random number of edges between 12 and 16, roughly matching what will be measured in the real data. This takes around a minute per dataset. We use four hidden layers with 50 neurons each and a Rectified Linear Unit (ReLU) activation function. The noise term $E_{i}$ is drawn independently for each sample from the default beta distribution Beta(E;2,5)(E;2,5). This is defined by
$\text{Beta}(E;\alpha,\beta)=\frac{E^{\alpha-1}(1-E)^{\beta-1}\Gamma(\alpha+\beta)}{\Gamma(\alpha)\Gamma(\beta)}$ (4)
where $\Gamma$ is the Gamma function. This skewed distribution produces a realistic amount of non-Gaussianity in the scatters, generating mock datasets with correlations visually similar to Fig. 1. We nevertheless check explicitly that the results are not sensitive to reasonable variations in the noise (see also Sec. 4.1).
We then refit each of these datasets with the FCIT algorithm for a range of truncation_limit and penalty_discount values. For each one we compute the precision (fraction of predicted edges that are correct), recall (fraction of true edges that were successfully recovered) and F1 score (harmonic mean of the precision and recall) of the PAG produced.
@@ 4.1 Mock data tests
The result is shown in Fig. 3. As penalty_discount increases the score penalises model complexity more strongly, leading to sparser graphs. This reduces false positives and thus tends to increase precision, but also causes some true edges to be missed, lowering recall. The F1 score peaks at penalty_discount $\sim 40-50$ at a value $\sim 0.9$, roughly indicating a 90 per cent success rate on each dataset. We adopt a value of 50 for the real data.
While the mock data matches the real data in the number of datapoints, nodes and (approximately, since this is a post-hoc statistic) causal edges, it does not match it in noise distribution, which is unknown and likely complex in the NSA. This may be expected to incur some systematic error in the optimal trunction_limit and penalty_discount when applied to the real data. However, the correlations induced in the mock data are qualitatively similar to those of Fig. 1, and we explicitly check that the results of Fig. 3 are not appreciably altered by reasonable variations in the neural network architecture or noise distribution parameters. Further, we investigate modest variations in trunction_limit, penalty_discount and the other hyperparameters in the FCIT algorithm and its testing and scoring methods directly on the real data, finding relatively robust results.
@@ 4.2 Real data
The result describes a combination of physical effects carrying information about galaxy evolution and observational and selection effects describing the way in which the data was obtained. As expected, redshift determines apparent size, which scales inversely with angular diameter distance. It also influences mass and absolute magnitude through Malmquist bias, the preferential detection of intrinsically brighter objects at higher distance (Eddington, 1914; Malmquist, 1922). Larger stellar mass leads to higher luminosity (ABSMAG) simply because more stars produce more light. Likewise, we expect more massive (or more luminous) galaxies to be larger: in SDSS the median half-light radius grows with mass as roughly $R\propto M^{0.4}$ for late types and even steeper ($\sim M^{0.55}$) for early types (Shen et al., 2003). Our finding ABSMAG $\rightarrow$ ELPETRO_TH50_R is consistent with these scaling laws (Courteau et al., 2007; Bernardi et al., 2014; Lange et al., 2015), but crucially exposes directional causal information. Late-type (disk) galaxies follow a shallow size–mass relation ($R_{\rm eff}\propto M_{*}^{0.22}$ for $M_{*}\gtrsim 3\times 10^{9}\,M_{\odot}$), whereas early types follow a steeper relation ($R_{\rm eff}\propto M_{*}^{0.75}$) (van der Wel et al., 2014). The arrow from luminosity to size therefore reflects the well-known mass–size relation in galaxy surveys (Shen et al., 2003; van der Wel et al., 2014), although it is interesting that it is luminosity rather than mass that appears to be the fundamental driver.
The less obvious edges, and the graph structure as a whole, may be interpreted as follows:
  • Size growth and disk assembly. The absence of a direct edge between recent star formation (log(B300)) and size indicates no evidence that short-timescale star formation independently sets galaxy size once stellar mass, luminosity, and morphology are accounted for. Nevertheless, the causal chain log(B300) $\rightarrow$ ABSMAG $\rightarrow$ ELPETRO_TH50_R remains physically meaningful, implying inside-out disk growth as a plausible physical interpretation rather than a directly inferred causal link (Muñoz-Mateos et al., 2011; Nelson et al., 2016). Observationally, galaxies with extended, actively star-forming outer disks exhibit larger blue-band radii and colour gradients consistent with gradual disk build-up at large radii (Wang et al., 2011). Overall, the PAG disfavours models in which instantaneous star formation efficiency alone drives size growth, instead supporting scenarios where size is primarily regulated by longer-timescale mass assembly and structural evolution, with star formation acting as a secondary or correlated process (Brook et al., 2012; Pilkington et al., 2012).
  • Merger-driven size evolution. The directed influence of luminosity (and implicitly mass) on size is also naturally interpreted in terms of merger-driven growth. Massive early-type galaxies are known to experience substantial size evolution through minor and major mergers at late times (van Dokkum et al., 2010; López-Sanjuan et al., 2012; Hilz et al., 2013). Since $z\sim 1$, massive ellipticals have undergone on average $\sim 0.9$ mergers, sufficient to explain $\sim 50$–75 per cent of their observed increase in half-light radius (López-Sanjuan et al., 2012). The recovered luminosity-to-size pathway is consistent with this hierarchical picture, in which mass assembly through mergers increases galaxy size more efficiently than stellar mass alone.
  • Mass, luminosity, and star formation. The edge log(MASS) $\rightarrow$ ABSMAG reflects the fundamental mass–luminosity relation. The additional edge log(B300) $\rightarrow$ ABSMAG captures the contribution of young stellar populations to optical luminosity, with recent star formation temporarily brightening galaxies. Together, these edges indicate that luminosity encodes both long-term mass assembly and short-term star formation activity. The partially directed connection between star formation and mass is consistent with the star-forming main sequence, ${\rm SFR}\propto M_{*}^{\alpha}$ with $\alpha\sim 0.6$ at low redshift (Whitaker et al., 2012; Speagle et al., 2014), while allowing for latent confounding by gas supply or environment (Peng et al., 2010; Tacconi et al., 2018).
  • Morphology as a mediator. The Sérsic index (SERSIC_N) plays a central role in the PAG, with a directed edge to size and uncertain connections to both stellar mass and star formation. As a proxy for bulge dominance, $n\simeq 1$ traces disk-dominated systems, while $n\simeq 4$ corresponds to bulge-dominated galaxies (Graham and Driver, 2005; Conselice, 2014). The edge SERSIC_N $\rightarrow$ ELPETRO_TH50_R indicates that morphology causally influences size at fixed luminosity, consistent with the fact that bulge-dominated systems are more compact. Here “causal” should be understood in the structural sense of the PAG: SERSIC_N acts as a proxy for the cumulative outcome of physical processes such as merging, disk instabilities, and quenching, rather than as a fundamental dynamical variable (Kormendy and Kennicutt, 2004). The circle endpoints on edges linking Sérsic index to mass and star formation rate indicate uncertainty in causal direction or the presence of latent variables, such as merger history or environment, that jointly affect morphology, mass growth, and quenching.
  • Uncertain edges and latent confounding. Several further edges appear with circle endpoints, particularly among morphology, mass, and star formation. These likely reflect the influence of unobserved confounders including halo mass, gas accretion history, feedback efficiency, and large-scale environment. Similarly, ambiguous links involving the apparent axis ratio (ELPETRO_BA) may arise from inclination-dependent dust attenuation, which affects apparent luminosity and measured structural parameters. In the PAG formalism, such circle endpoints explicitly encode the limits of identifiability given the conditional independence structure of the available variables and sample size.
  • Synthesis with galaxy-formation theory. Overall, the PAG supports a picture in which stellar mass assembly underlies galaxy evolution, setting luminosity and, through it, galaxy size (Naab and Ostriker, 2017). Morphology acts as a key mediator, reflecting the cumulative impact of mergers and internal dynamical processes, while recent star formation modulates luminosity without independently determining size. The recovered structure is consistent with observational evidence that, at fixed mass, late-type galaxies are larger than early types and that galaxy sizes evolve as $R_{\rm eff}\propto(1+z)^{-0.75}$ for disks and $\sim(1+z)^{-1.48}$ for ellipticals (van der Wel et al., 2014). The PAG therefore captures well-established empirical regularities and distinguishing between plausible physical pathways that would otherwise be empirically degenerate, while clearly delineating where the data is not sufficient to support stronger causal claims. This is highly promising for the future, where more sophisticated applications of the methodology to data of higher quality and quantity may yield crucial insight into the physical mechanisms of galaxy formation.
@@ 5 Discussion and Conclusion
To illustrate the approach we have applied causal discovery to low-redshift galaxy data from the NSA, adapting a hybrid constraint-and-score algorithm—FCIT—to meet the demands of astrophysical data (large datasets, highly nonlinear correlations and presence of confounders). After testing and calibrating the method on NSA-like mock data (establishing $\sim$90 per cent accuracy) we applied it to the real data to find the PAG in Fig. 4. This supports a hierarchical, morphology- and mass-driven framework of galaxy evolution while indicating the complexities involved in the physical mechanisms at play. It also highlights the vital importance of observational causal discovery methods, since intervention is impossible in astrophysics. Such methods are sometimes undervalued in favour of intervention-based methods involving randomised control trials.
@@
General and efficient causal discovery methods remain under active development. Future algorithms might be able to resolve circle endpoints to distinguish models within a PAG equivalence class by utilising other types of information (e.g. nonlinearity, non-Gaussianity or heteroskedasticity), as has already been done when assuming causal sufficiency in the linear, non-Gaussian regime (e.g. Hyvärinen and Zhang 2010; Shimizu et al. 2011). One could even do Bayesian model comparison between competing simulations or theories based on their causal structures they predict (Dhir et al., 2024; Jin et al., 2025b), directly demonstrating the gain in constraining power afforded by causal discovery. Taken together, these steps would provide a more complete and reliable causal map, ultimately allowing causal discovery to move beyond recovering broad backbones to testing detailed models of how galaxies grow and evolve.
@@
The Nasa Sloan Atlas is publicly available at https://www.sdss4.org/dr17/manga/manga-target-selection/nsa/. The code used in this paper is publicly available on Github .
@@ Acknowledgements
We thank David Bacon, Deaglan Bartlett, Robin Evans, Pedro Ferreira, Sebastian von Hausegger, Matt Jarvis, Richard Stiskalek, Simon White and Tariq Yasin for useful discussions. HD is supported by a Royal Society University Research Fellowship (grant no. 211046).